OpenStudy (idealist10):

A spring-mass system has natural frequency 7sqrt(10) rad/s. The natural length of the spring is 0.7 m. Find the length of the spring when the mass is in equilibrium.

OpenStudy (idealist10):

\[y''+\frac{ k }{ m }y=0\]

OpenStudy (idealist10):

@freckles @jim_thompson5910 @Kainui @mathstudent55 @SolomonZelman @Zarkon

OpenStudy (solomonzelman):

Undamped free case:)

OpenStudy (solomonzelman):

Let \(\color{#000000}{ \displaystyle \omega:=\sqrt{\frac{k}{m}}}\) \(\color{#000000}{ \displaystyle y''+\frac{k}{m}y=0\quad \Rightarrow \quad y''+\omega^2y=0}\) \(\color{#000000}{ \displaystyle r^2+\omega^2=0}\) \(\color{#000000}{ \displaystyle r=\pm\omega{\tiny~} i}\) and then the "prove" for a simple harmonica motion formula (by letting \(c_1=A\sin\phi\) and \(c_2=A\cos\phi\), and re-writing) but, you don't want to discuver formulas, just solve it, right?

OpenStudy (solomonzelman):

You can write the auxiliary equation \(\color{#000000}{ \displaystyle r^2+\frac{k}{m}=0}\) ===> \(\color{#000000}{ \displaystyle r=\pm i\sqrt{\frac{k}{m}}}\) what is exactly the problem tho?

OpenStudy (idealist10):

So suppose that